# FedFACT: A Provable Framework for Controllable Group-Fairness Calibration in Federated Learning

### v1

@@ B.3 Proof of Proposition 3 and Further Exploration
This lemma is commonly used in the design of multiclass loss functions [69, 52, 54].
@@ B.3.1 Proof of Proposition 3
$\mathbb{E}$ [${}_{(x,y,a,k)\sim(X,Y,A,K)}[\ell_{k}(y,\mathbf{s}(x),a)]=-\mathbb{E}_{X,Y,A,K}\Bigg{[}\sum_{i=1}^{m}\overline{\mathbf{M}}^{\lambda,\mu}_{Y,i}(A,K)\log\frac{\exp([\mathbf{s}(X)]_{i})}{\sum_{j=1}^{m}\exp([\mathbf{s}(X)]_{j})}\Bigg{]}$
@@
It presents that, for all $i\in[m]$, since $\sum_{i\in[m]}\eta(x,a,k)=1$,
@@ B.3.2 Exploration of Calibrated Loss for Unified Bayes-Optimal Classifier
$\mathbb{E}$ [${}_{(x,y,a,k)\sim(X,Y,A,K)}[\ell_{k}(y,\mathbf{s}(x),a)]=-\mathbb{E}_{X,Y,A,K}\Bigg{[}\sum_{i=1}^{m}\overline{\mathbf{M}}^{\lambda,\mu}_{Y,i}(A,K)\log\frac{\exp([\mathbf{s}(X)]_{i})}{\sum_{j=1}^{m}\exp([\mathbf{s}(X)]_{j})}\Bigg{]}$
@@
Assumption 2 and 3 are standard in the convergence analysis of federated model [47, 45, 66]. Now we present Theorem 7 and Theorem 8, which together constitute an extended form of Theorem 4.
@@
Suppose that personalized models achieve a $\rho_{t}$-approximate optimal response at iteration $t$, namely ^$\widehat{\mathcal{L}}(\mathbf{h}^{t},\lambda^{t},\mu^{t})\leq\min_{\mathbf{h}}\widehat{\mathcal{L}}(\mathbf{h},\lambda^{t},\mu^{t})+\rho_{t}$, denoting $\bar{\rho}=\sum_{t=1}^{T}\rho_{t}/T$, then the sequences of model and bounded dual parameters comprise an approximate mixed Nash equilibrium:
@@ B.4.1 Proof of Theorem 7
Here, the update can be viewed as exponentiated gradient descent on the normalized weight vector $\mathbf{w^{t}_{k}}=(w^{t}_{k,1},w^{t}_{k,2})\in\Delta_{2}$, and ${w^{t}_{k,i}}\mathbb{P}opto\exp{(-\eta_{w}z^{k}_{t,i})},i=1,2$, where $z^{k}_{t,1}=L^{t}_{k}(f(x;\theta^{t})),z^{k}_{t,2}=L^{t}_{k}(f(x;\phi^{t}_{k}))$. A well-known regret bound in online learning [63, 64] shows that, for any $\mathbf{u}=(u_{1},u_{2})\in\Delta_{2}$,
@@ B.4.2 Proof of Theorem 8
[63] Let $f^{1},f^{2},\ldots:\Lambda\rightarrow\mathbb{R}$ be a sequence of convex functions that we wish to minimize on a compact convex set $\Lambda$. Define the bound of the convex set $B_{d}\geq\max_{\lambda\in\Lambda}\|\Lambda\|_{2}$, and $B_{G}\geq\left\|{\nabla}f^{t}\left(\lambda^{t}\right)\right\|_{2}$ is a uniform upper bound on the norms of the subgradients. Suppose that we perform $T$ iterations of the following update, starting from $\lambda^{(1)}=\operatorname{argmin}_{\lambda\in\Lambda}\|\lambda\|_{1}$:
@@
Proof of Theorem 8. Consider the empirical form of the Lagrangian function ^$\widehat{\mathcal{L}}(\mathbf{h},\lambda,\mu)$,
$\widehat{\mathcal{L}}($ $\mathbf{h},\lambda,\mu)=\widehat{\mathcal{R}}(\mathbf{h})+(\lambda^{(1)}-\lambda^{(2)})^{\top}(\widehat{\mathscr{D}}^{g}(\mathbf{h})-\xi^{g})+\sum_{k=1}^{N}(\mu^{1,k}-\mu^{2,k})^{\top}(\widehat{\mathscr{D}}^{l,k}(\mathbf{h})-\xi^{l,k}),$
@@
where $\widehat{\mathbf{M}}^{\lambda,\mu}(a,k):=\mathbf{I}-\frac{1}{\hat{p}_{a,k}}\Big{[}\sum_{u_{g}\in\mathcal{U}_{g}}(\lambda^{(1)}_{u_{g}}-\lambda^{(2)}_{u_{g}})\widehat{\mathbf{D}}^{a,k}_{u_{g}}-\sum_{u_{l,k}\in\mathcal{U}_{l,k}}(\mu^{(1)}_{k,u_{l,k}}-\mu^{(2)}_{k,u_{l,k}})\widehat{\mathbf{D}}^{a,k}_{u_{l,k}}\Big{]}$. It is clear that the inner problem is linear to classifiers in the empirical case.
@@
Let $\overline{\mathbf{h}}:=\frac{1}{T}\sum_{t=1}^{T}\mathbf{h}^{t}$ with $\overline{h}_{k}:=\frac{1}{T}\sum_{t=1}^{T}{h}^{t}_{k},k\in[N]$, and let $\bar{\lambda}:=\frac{1}{T}\sum_{t=1}^{T}\lambda^{t},\bar{\mu}:=\frac{1}{T}\sum_{t=1}^{T}\mu^{t}$ denote the point-wise average of dual parameters. Therefore, due to the linearity of empirical Lagrange function to classifiers and dual parameters, (25) can be formulated as
@@ B.5 Generalization Error For In-processing Algorithm
We begin by introducing some notations and simplifications, which are commonly employed in generalization analyses of FL [35, 61]. Without loss of generalization, let $n=n_{1}=\cdots=n_{k}$ present the sample number in local datasets. For any class =$\mathcal{H}=\{h:\mathcal{X}\to[m]\}$, denote $\mathcal{H}_{y}=\left\{\mathbb{I}{\{h(x)=y\}}:h\in\mathcal{H}\right\}$ and the maximal Vapnik-Chervonenkis dimension [62], $VC(\mathcal{H}):=\max_{y\in[m]}VC(\mathcal{H}_{y})$.
@@
If classifiers $\overline{\mathbf{h}}=(\bar{h}_{1},\dots,\bar{h}_{k})$ with dual parameters $(\bar{\lambda},\bar{\mu})$ form a $\epsilon$-saddle point of empirical Lagrangian ^$\widehat{\mathcal{L}}(\mathbf{h},\lambda,\mu)$, and an optimal solution $\mathbf{h}^{*}\in\mathcal{H}$ satisfies both global and local fairness constraints, denoting $\nu(n,\mathcal{H},\delta)=2\sqrt{\frac{2VC(\mathcal{H})\log(n+1)}{n}}+\sqrt{\frac{2\log(m^{2}N/\delta))}{n}}$, $B_{g}=\max_{a\in\mathcal{A},k\in[N]}\|\mathbf{D}^{a,k}_{u_{g}}\|_{1},\Omega_{n}^{g}=\max_{a\in\mathcal{A},k\in[N]}\|\mathbf{D}^{a,k}_{u_{g}}-\widehat{\mathbf{D}}^{a,k}_{u_{g}}\|_{\infty}$, $\Omega^{p}_{n}:=\sum_{k=1}^{N}|p_{k}-\hat{p}_{k}|$, and $B_{l,k}=\max_{a\in\mathcal{A}}\|\mathbf{D}^{a,k}_{u_{l,k}}\|_{1},\Omega_{n}^{l,k}=\max_{a\in\mathcal{A}}\|\mathbf{D}^{a,k}_{u_{l,k}}-\widehat{\mathbf{D}}^{a,k}_{u_{l,k}}\|_{\infty},k\in[N]$, then with probability at least $1-\delta$,
@@
$|\mathscr{D}^{l,k}(\overline{\mathbf{h}})|\leq\xi^{l,k}+\nu(n,\mathcal{H},N\delta/|\mathcal{A}||\mathcal{U}_{l,k}|)|\mathcal{A}|B_{l,k}+\Omega_{n}^{l,k}+\frac{1+2\epsilon}{B_{d}}$
@@
where :=$\mathcal{F}_{\mathcal{H}}:=\{f(x)=\sum_{j=1}^{N}\alpha_{j}h_{j}(x):\alpha\in\Delta_{N},h_{j}\in\mathcal{H},j\in[m]\}$.
Proof of Lemma B.3.  Let $\ell_{i,j}(x,y;h)=\mathbb{I}(y=i\wedge h(x)=j)$. Then we have $\mathbf{C}_{i,j}(h)=\mathbb{E}[\ell_{i,j}(x,y;h)]$ and $\widehat{\mathbf{C}}_{i,j}(h)=\frac{1}{n}\sum_{i=1}^{n}\ell_{i,j}(x_{i},y_{i};h)$. Hence, according to the classical result with respect to cost sensitive binary classification [7], with probability at least $1-\delta$,
@@ B.5.1 Proof of Theorem 9
^$\widehat{\mathcal{L}}\left(\overline{\mathbf{h}},\bar{\lambda},\bar{\mu}\right)\leq\widehat{\mathcal{L}}\left(\mathbf{h},\bar{\lambda},\bar{\mu}\right)+\epsilon,\qquad$ $\forall\ \mathbf{h}\in\mathcal{H},$ (27)
^$\widehat{\mathcal{L}}\left(\overline{\mathbf{h}},\bar{\lambda},\bar{\mu}\right)\geq\widehat{\mathcal{L}}\left(\overline{\mathbf{h}},\lambda,\mu\right)-\epsilon,\qquad$ $\forall\ \|\lambda\|_{1},\|\mu\|_{1}\leq B_{d}.$ (28)
@@
Since $\|\widehat{\mathbf{C}}^{a,k}(h_{k})\|_{1}=1$, taking union bound again, denoting $\nu(n,\mathcal{H},\delta)=2\sqrt{\frac{2VC(\mathcal{H})\log(n+1)}{n}}+\sqrt{\frac{2\log(m^{2}N/\delta))}{n}}$, it turns out that with probability at least $1-\delta$,
@@
^$\widehat{\mathcal{L}}(\overline{\mathbf{h}},\bar{\lambda},\bar{\mu})-\widehat{R}(\overline{\mathbf{h}})\leq\widehat{\mathcal{L}}({\mathbf{h}},\bar{\lambda},\bar{\mu})-\widehat{R}(\overline{\mathbf{h}})+\epsilon\leq\widehat{R}({\mathbf{h}})-\widehat{R}(\overline{\mathbf{h}})+\epsilon.$ (32)
@@
For local fairness constraints, $|\mathscr{D}_{l,k}|\leq\xi^{l,k}$, following the similar proof procedures as local fairness constraints, we have that
$|{\mathscr{D}}^{l,k}(\overline{\mathbf{h}})|-\xi^{l,k}$ $\leq\nu(n,\mathcal{H},N\delta/|\mathcal{A}||\mathcal{U}_{l,k}|)|\mathcal{A}|B_{l,k}+\Omega^{l,k}_{n}+\frac{1+2\epsilon}{B_{d}}.$ (37)
@@
^$\widehat{\mathcal{R}}(\overline{\mathbf{h}})-\widehat{\mathcal{R}}(\mathbf{h}^{*})\leq\widehat{\mathcal{L}}(\overline{\mathbf{h}},\mathbf{0},\mathbf{0})-\widehat{\mathcal{L}}(\mathbf{h}^{*},\bar{\lambda},\bar{\mu})\leq\widehat{\mathcal{L}}\left(\overline{\mathbf{h}},\bar{\lambda},\bar{\mu}\right)+\epsilon-\widehat{\mathcal{L}}\left(\overline{\mathbf{h}},\bar{\lambda},\bar{\mu}\right)+\epsilon=2\epsilon.$ (39)
Since we have ^$\widehat{\mathcal{R}}(\mathbf{h})=1-\sum_{k=1}^{N}\hat{p}_{k}\langle\mathbf{I},\mathbf{C}^{k}(h_{k})\rangle$, it presents that
^$\widehat{\mathcal{R}}(h)-{\mathcal{R}}(h)$ $=\sum_{k=1}^{N}\langle\mathbf{I},p_{k}\mathbf{C}^{k}(h_{k})-\hat{p}_{k}\widehat{\mathbf{C}}^{k}(h_{k})\rangle$
@@
^$\widehat{\mathcal{R}}(h)-{\mathcal{R}}(h)$ $\leq N\max_{k\in[N]}|p_{k}-\hat{p}_{k}|+\sum_{k=1}^{N}p_{k}m\|\mathbf{C}^{k}(h_{k})-\widehat{\mathbf{C}}^{k}(h_{k})\|_{\infty}$

### v2

@@ B.3 Proof of Proposition 3 and Further Exploration
This lemma is commonly used in the design of multiclass loss functions [lemma-multiclass-loss-JMLR:v17:14-294, lemma-multiclass-loss-menon2021longtail, lemma-multiclass-loss-over-parameterized].
@@ B.3.1 Proof of Proposition 3
$\mathbb{E}$ $_{(x,y,a,k)\sim(X,Y,A,K)}[\ell_{k}(y,\mathbf{s}(x),a)]=-\mathbb{E}_{X,Y,A,K}\Bigg[\sum_{i=1}^{m}\overline{\mathbf{M}}^{\lambda,\mu}_{Y,i}(A,K)\log\frac{\exp([\mathbf{s}(X)]_{i})}{\sum_{j=1}^{m}\exp([\mathbf{s}(X)]_{j})}\Bigg]$
@@
It presents that, for all $i\in[m]$, since $\sum_{i\in[m]}\eta_{i}(x,a,k)=1$,
@@ B.3.2 Exploration of Calibrated Loss for Unified Bayes-Optimal Classifier
$\mathbb{E}$ $_{(x,y,a,k)\sim(X,Y,A,K)}[\ell_{k}(y,\mathbf{s}(x),a)]=-\mathbb{E}_{X,Y,A,K}\Bigg[\sum_{i=1}^{m}\overline{\mathbf{M}}^{\lambda,\mu}_{Y,i}(A,K)\log\frac{\exp([\mathbf{s}(X)]_{i})}{\sum_{j=1}^{m}\exp([\mathbf{s}(X)]_{j})}\Bigg]$
@@
Assumption 2 and 3 are standard in the convergence analysis of federated model [convergence-of-fedavg-li2019convergence, ditto-li2021ditto, fednest-convergence-pmlr-v162-tarzanagh22a]. Now we present Theorem 7 and Theorem 8, which together constitute an extended form of Theorem 4.
@@
Suppose that personalized models achieve a $\rho_{t}$-approximate optimal response at iteration $t$, namely $\widehat{\mathcal{L}}(\mathbf{h}^{t},\lambda^{t},\mu^{t})\leq\min_{\mathbf{h}}\widehat{\mathcal{L}}(\mathbf{h},\lambda^{t},\mu^{t})+\rho_{t}$, denoting $\bar{\rho}=\sum_{t=1}^{T}\rho_{t}/T$, then the sequences of model and bounded dual parameters comprise an approximate mixed Nash equilibrium:
@@ B.4.1 Proof of Theorem 7
Here, the update can be viewed as exponentiated gradient descent on the normalized weight vector $\mathbf{w^{t}_{k}}=(w^{t}_{k,1},w^{t}_{k,2})\in\Delta_{2}$, and ${w^{t}_{k,i}}\propto\exp{(-\eta_{w}z^{k}_{t,i})},i=1,2$, where $z^{k}_{t,1}=L^{t}_{k}(f(x;\theta^{t})),z^{k}_{t,2}=L^{t}_{k}(f(x;\phi^{t}_{k}))$. A well-known regret bound in online learning [online-learning-textbook-shalev2012online, online-learning-regret-proof-steinhardt2014adaptivity] shows that, for any $\mathbf{u}=(u_{1},u_{2})\in\Delta_{2}$,
@@ B.4.2 Proof of Theorem 8
[online-learning-textbook-shalev2012online] Let $f^{1},f^{2},\ldots:\Lambda\rightarrow\mathbb{R}$ be a sequence of convex functions that we wish to minimize on a compact convex set $\Lambda$. Define the bound of the convex set $B_{d}\geq\max_{\lambda\in\Lambda}\|\Lambda\|_{2}$, and $B_{G}\geq\left\|{\nabla}f^{t}\left(\lambda^{t}\right)\right\|_{2}$ is a uniform upper bound on the norms of the subgradients. Suppose that we perform $T$ iterations of the following update, starting from $\lambda^{(1)}=\operatorname{argmin}_{\lambda\in\Lambda}\|\lambda\|_{1}$:
@@
Proof of Theorem 8. Consider the empirical form of the Lagrangian function $\widehat{\mathcal{L}}(\mathbf{h},\lambda,\mu)$,
^($\widehat{\mathcal{L}}$( $\mathbf{h},\lambda,\mu)=\widehat{\mathcal{R}}(\mathbf{h})+(\lambda^{(1)}-\lambda^{(2)})^{\top}(\widehat{\mathscr{D}}^{g}(\mathbf{h})-\xi^{g})+\sum_{k=1}^{N}(\mu^{1,k}-\mu^{2,k})^{\top}(\widehat{\mathscr{D}}^{k}(\mathbf{h})-\xi^{k}),$
@@
where $\widehat{\mathbf{M}}^{\lambda,\mu}(a,k):=\mathbf{I}-\frac{1}{\hat{p}_{a,k}}\Big[\sum_{u_{g}\in\mathcal{U}_{g}}(\lambda^{(1)}_{u_{g}}-\lambda^{(2)}_{u_{g}})\widehat{\mathbf{D}}^{a,k}_{u_{g}}-\sum_{u_{k}\in\mathcal{U}_{k}}(\mu^{(1)}_{k,u_{k}}-\mu^{(2)}_{k,u_{k}})\widehat{\mathbf{D}}^{a,k}_{u_{k}}\Big]$. It is clear that the inner problem is linear to classifiers in the empirical case.
@@
Let $\overline{\mathbf{h}}:=\frac{1}{T}\sum_{t=1}^{T}\mathbf{h}^{t}$ with $\overline{h}_{k}:=\frac{1}{T}\sum_{t=1}^{T}{h}^{t}_{k},k\in[N]$, and let $\bar{\lambda}:=\frac{1}{T}\sum_{t=1}^{T}\lambda^{t},\bar{\mu}:=\frac{1}{T}\sum_{t=1}^{T}\mu^{t}$ denote the point-wise average of dual parameters. Therefore, due to the linearity of the empirical Lagrange function to classifiers and dual parameters, (25) can be formulated as
@@ B.5 Generalization Error For In-processing Algorithm
We begin by introducing some notations and simplifications, which are commonly employed in generalization analyses of FL [federated-generalization-hu2023generalization, federated-PAC-generalization-sefidgaran2024lessons]. Without loss of generalization, let $n=n_{1}=\cdots=n_{k}$ present the sample number in local datasets. For any class $\mathcal{H}=\{h:\mathcal{X}\to[m]\}$, denote $\mathcal{H}_{y}=\left\{\mathbb{I}{\{h(x)=y\}}:h\in\mathcal{H}\right\}$ and the maximal Vapnik-Chervonenkis dimension [Understanding-Machine-Learning-textbook], $VC(\mathcal{H}):=\max_{y\in[m]}VC(\mathcal{H}_{y})$.
@@
If classifiers $\overline{\mathbf{h}}=(\bar{h}_{1},\dots,\bar{h}_{k})$ with dual parameters $(\bar{\lambda},\bar{\mu})$ form a $\epsilon$-saddle point of empirical Lagrangian $\widehat{\mathcal{L}}(\mathbf{h},\lambda,\mu)$, and an optimal solution $\mathbf{h}^{*}\in\mathcal{H}$ satisfies both global and local fairness constraints, denoting $\nu(n,\mathcal{H},\delta)=2\sqrt{\frac{2VC(\mathcal{H})\log(n+1)}{n}}+\sqrt{\frac{2\log(m^{2}N/\delta))}{n}}$, $B_{g}=\max_{a\in\mathcal{A},k\in[N]}\|\mathbf{D}^{a,k}_{u_{g}}\|_{1},\Omega_{n}^{g}=\max_{a\in\mathcal{A},k\in[N]}\|\mathbf{D}^{a,k}_{u_{g}}-\widehat{\mathbf{D}}^{a,k}_{u_{g}}\|_{\infty}$, $\Omega^{p}_{n}:=\sum_{k=1}^{N}|p_{k}-\hat{p}_{k}|$, and $B_{k}=\max_{a\in\mathcal{A}}\|\mathbf{D}^{a,k}_{u_{k}}\|_{1},\Omega_{n}^{k}=\max_{a\in\mathcal{A}}\|\mathbf{D}^{a,k}_{u_{k}}-\widehat{\mathbf{D}}^{a,k}_{u_{k}}\|_{\infty},k\in[N]$, then with probability at least $1-\delta$,
@@
$|\mathscr{D}^{k}(\overline{\mathbf{h}})|\leq\xi^{k}+\nu(n,\mathcal{H},N\delta/|\mathcal{A}||\mathcal{U}_{k}|)|\mathcal{A}|B_{k}+\Omega_{n}^{k}+\frac{1+2\epsilon}{B_{d}}$
@@
where $\mathcal{F}_{\mathcal{H}}:=\{f(x)=\sum_{j=1}^{N}\alpha_{j}h_{j}(x):\alpha\in\Delta_{N},h_{j}\in\mathcal{H},j\in[m]\}$.
Proof of Lemma B.3. Let $\ell_{i,j}(x,y;h)=\mathbb{I}(y=i\wedge h(x)=j)$. Then we have $\mathbf{C}_{i,j}(h)=\mathbb{E}[\ell_{i,j}(x,y;h)]$ and $\widehat{\mathbf{C}}_{i,j}(h)=\frac{1}{n}\sum_{i=1}^{n}\ell_{i,j}(x_{i},y_{i};h)$. Hence, according to the classical result with respect to cost sensitive binary classification [binary-classification-generalization-bound-boucheron2005theory], with probability at least $1-\delta$,
@@ B.5.1 Proof of Theorem 9
$\widehat{\mathcal{L}}\left(\overline{\mathbf{h}},\bar{\lambda},\bar{\mu}\right)\leq\widehat{\mathcal{L}}\left(\mathbf{h},\bar{\lambda},\bar{\mu}\right)+\epsilon,\qquad$ $\forall\ \mathbf{h}\in\mathcal{H},$ (27)
$\widehat{\mathcal{L}}\left(\overline{\mathbf{h}},\bar{\lambda},\bar{\mu}\right)\geq\widehat{\mathcal{L}}\left(\overline{\mathbf{h}},\lambda,\mu\right)-\epsilon,\qquad$ $\forall\ \|\lambda\|_{1},\|\mu\|_{1}\leq B_{d}.$ (28)
@@
Since $\|\widehat{\mathbf{C}}^{a,k}(h_{k})\|_{1}=1$, taking the union bound again, denoting $\nu(n,\mathcal{H},\delta)=2\sqrt{\frac{2VC(\mathcal{H})\log(n+1)}{n}}+\sqrt{\frac{2\log(m^{2}N/\delta))}{n}}$, it turns out that with probability at least $1-\delta$,
@@
$\widehat{\mathcal{L}}(\overline{\mathbf{h}},\bar{\lambda},\bar{\mu})-\widehat{R}(\overline{\mathbf{h}})\leq\widehat{\mathcal{L}}({\mathbf{h}},\bar{\lambda},\bar{\mu})-\widehat{R}(\overline{\mathbf{h}})+\epsilon\leq\widehat{R}({\mathbf{h}})-\widehat{R}(\overline{\mathbf{h}})+\epsilon.$ (32)
@@
For local fairness constraints, $|\mathscr{D}_{k}|\leq\xi^{k}$, following the similar proof procedures as local fairness constraints, we have that
$|{\mathscr{D}}^{k}(\overline{\mathbf{h}})|-\xi^{k}$ $\leq\nu(n,\mathcal{H},N\delta/|\mathcal{A}||\mathcal{U}_{k}|)|\mathcal{A}|B_{k}+\Omega^{k}_{n}+\frac{1+2\epsilon}{B_{d}}.$ (37)
@@
$\widehat{\mathcal{R}}(\overline{\mathbf{h}})-\widehat{\mathcal{R}}(\mathbf{h}^{*})\leq\widehat{\mathcal{L}}(\overline{\mathbf{h}},\mathbf{0},\mathbf{0})-\widehat{\mathcal{L}}(\mathbf{h}^{*},\bar{\lambda},\bar{\mu})\leq\widehat{\mathcal{L}}\left(\overline{\mathbf{h}},\bar{\lambda},\bar{\mu}\right)+\epsilon-\widehat{\mathcal{L}}\left(\overline{\mathbf{h}},\bar{\lambda},\bar{\mu}\right)+\epsilon=2\epsilon.$ (39)
Since we have $\widehat{\mathcal{R}}(\mathbf{h})=1-\sum_{k=1}^{N}\hat{p}_{k}\langle\mathbf{I},\mathbf{C}^{k}(h_{k})\rangle$, it presents that
$\widehat{\mathcal{R}}(h)-{\mathcal{R}}(h)$ $=\sum_{k=1}^{N}\langle\mathbf{I},p_{k}\mathbf{C}^{k}(h_{k})-\hat{p}_{k}\widehat{\mathbf{C}}^{k}(h_{k})\rangle$
@@
$\widehat{\mathcal{R}}(h)-{\mathcal{R}}(h)$ $\leq N\max_{k\in[N]}|p_{k}-\hat{p}_{k}|+\sum_{k=1}^{N}p_{k}m\|\mathbf{C}^{k}(h_{k})-\widehat{\mathbf{C}}^{k}(h_{k})\|_{\infty}$